OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through (4, 2) and (-2, -3)?

OpenStudy (austinl):

First you need to find the slope, \(\large{m=\dfrac{\color{red}{y_2}-\color{blue}{y_1}}{\color{green}{x_2}-\color{orange}{x_1}}}\) With the points, \((\color{orange}{-2},\color{blue}{-3})\) \((\color{green}{4},\color{red}{2})\)

OpenStudy (austinl):

I have to go now, I am sure someone else can take over from here! Good luck!

OpenStudy (anonymous):

slope m = 2 - (-3) / 4 - (-2) = 5/6 y - 2 = 5/6(x -4) y = 5/6x - 4/3

OpenStudy (austinl):

@s01025881 we try not to just give answers here. Guidance toward answers is what we do.